OpenStudy (anonymous):

How many solutions does this system of equations have? 5x - 4y = 11 4y = 5x - 9

OpenStudy (anonymous):

How many solutions does this system of equations have? 5x - 4y = 11=x=4n+3 y=5n+1

OpenStudy (anonymous):

None, because if you simplify it, it becomes: 5x-4y=11 5x-4y=9 not possible

OpenStudy (anonymous):

MathFan12 is exactly correct. This is called an inconsistent system; the contradiction (eleven and nine are clearly not the same) says it has no solution.

OpenStudy (anonymous):

Thank You!